Refer to the following partial ANOVA results from EXCEL.

source of variation ss df MS F-sta
among-group 172.740 43.185
within-group 508.972
total

total # of observation: 58

e. the critical value , F x at x=0.05 (3 decimals)

f. at x= 0.05 your conclusion for testing the null hypothesis is that all the groups have the same mean is:
Answer #1

total observation (n) = 58

the necessary calculation table be:-

source of variation ss df MS= ss/df F-statistic
among-group 172.740

(172.740/43.185)

= 4

43.185

=(43.185 / 9.603)

= 4.497

within-group 508.972 (57-4) = 53

(508.972/53)

=9.603

total n-1 = (58-1) = 57

e).number of factors (r) = 4+1 = 5

[ because df among group = r-1 = 4]

The degrees of freedom for the numerator are r – 1 = (5-1) = 4

The degrees of freedom for the denominator are (nr) = ( 58 - 5) = 53

the critical value be:-

[ from f distribution table, for df1 = 4 , df2 = 53 , alpha = 0.05 ]

f).decision:-

so, we reject the null hypothesis.

hence we conclude that,

there is not enough evidence to claim that all the groups have the same mean at 0.05 level of significance.

*** if you face any trouble to understand the answer to the problem please mention it in the comment box.if you are satisfied, please give me a LIKE if possible.